A heat engine is used to drive a heat pump. The heat transferred from the heat engine and from the heat pump is used to heat the water circulating through the radiator of a building. The efficiency of the heat engine is 37% and the cop of the heat pump is 6. Evaluate the ratio of the heat transfer to the circulating water to the heat transfer to the heat engine

Grade:12th pass

1 Answers

Arun
5 years ago

Let Q1 be heat entering the engine and Q2 and Q3 be the heat transfer to water.

To find: (Q2+Q3)/Q1 = k

Given: η = 0.37; COP =6

Solution: W = ηQ1 = 0.37Q1

Q2 = (1-η) Q1 = 0.63 Q1

Q3 = COP x W = 6 x 0.37 x Q1 = 2.22 Q1

K = (0.63 Q1 + 2.22 Q1)/Q1 = 2.85

Regards

Arun